OpenStudy (cookie_2046):

A. Definition of Subtraction B. Additive inverse Property C. Additive Identity Property D. Opposite of a sum Property

OpenStudy (cookie_2046):

My question is on the attachment

OpenStudy (zzr0ck3r):

Any ideas?

OpenStudy (anonymous):

The additive inverse of a number x is another number, y, as long as the sum of x + y equals zero. In other words, the additive inverse of x is equal and opposite in sign to it (i.e., y = -x or vice versa). In mathematics the additive identity of a set which is equipped with the operation of addition is an element which, when added to any element x in the set, yields x. The property refers to how the opposite of a sum of real numbers is equal to the sum of the real numbers' opposites. The property written out is -(a+b)=(-a)+(-b). And of course, subtraction is just 3-1 = 2 and etcetc. Do these definitions help you out?

OpenStudy (zzr0ck3r):

You are missing a crucial one @rheaanderson341 Subtraction is defined as a-b=a+(-b)

OpenStudy (anonymous):

Oh wait. I never saw it as that! @zzr0ck3r . Guess we both learn something today Cookie!

OpenStudy (mdilley):

B